OpenStudy (anonymous):

solve for x using log 2^x+2^-x=3

OpenStudy (lgbasallote):

\[2^x + 2^{-x} = 3\] i remember seeing a problem like this before.. you know how to do this right @KingGeorge

OpenStudy (kinggeorge):

Ah one of these. First thing to do, is convert it to a quadratic.

OpenStudy (kinggeorge):

\[2^x+2^{-x}=3 \implies (2^x)^2-3\cdot2^x+1=0\]Note that this is a quadratic in \(2^x\). I don't believe this factors nicely, so now we use the quadratic formula.

OpenStudy (kinggeorge):

\[a=1\]\[b=-3\]\[c=1\]\[2^x=\frac{3\pm\sqrt{9-4}}{2}\]\[2^x=\frac{3\pm\sqrt{5}}{2}\]From here, look at \[2^x=\frac{3+\sqrt{5}}{2}\]\[2^x=\frac{3-\sqrt5}{2}\]and take log base 2 of both sides.

OpenStudy (anonymous):

thanks! how did you convert it to (2 x ) 2 −3⋅2 x +1=0 though?

OpenStudy (kinggeorge):

I multiplied by \(2^x\). So \[2^x\left(2^x+2^{-x}\right)=2^x\cdot3\]And then distribute and subtract \(3\cdot2^x\) from both sides.